The Bata Company in Islamabad reports the following statistics on annual basis.                                                                                           

Net Operating Income

RS: 30, 000

Average Operating Assests

RS: 200, 000

Sales

RS: 500, 000

Operating Expenses

RS: 470, 000

 

(a)  Calculate the Bata Company Return on Investment (ROI).

(b) Suppose that Bata’s manager invests in a RS. 30, 000 piece of equipment that increases sales by RS. 35,  

      000, while increasing operating expenses by RS. 15, 000. This increase alters the above mentioned 

      statistics. Calculate the new ROI.
Answer #1


(a): To calculate the ROI, we need to use a formula of ROI, which says the profit earned on an investment is divided by the cost of that investment to calculate return on investment (ROI).


So, to calculate the ROI for this scenario,

The total profit earned is = Net Operating Income + Average operating Assets + Sales / Operating Expenses


ROI = 30,000 + 200, 000 + 500, 000 / 470, 000

ROI = 7,30,000 / 470, 000

ROI = 1.55319148936


Now in order to calculate the new ROI, we need to consider new values


The total profit earned is = Net Operating Income + Average operating Assets + Sales / Operating Expenses


New ROI = 35,000 + 200, 000 + 500, 000 / 470,000 + 15,000

New ROI = 735000 / 485000

New ROI = 1.51546391753


So, the old ROI was 1.55319148936 which is 155.31% and New ROI is 1.51546391753 which is 151.54%


answered by: Allen